OpenStudy (anonymous):

does anyone know how to solve tho using matrices? {2x+5y=14 {x-y=7/2

hero (hero):

Do you have to use matrices to solve?

OpenStudy (anonymous):

yes

hero (hero):

[2 5 14] [1 -1 3.5]

OpenStudy (anonymous):

Now what you want it to get the diagonal from left to right all 1's and the rest 0's. This takes a series of row operations

OpenStudy (anonymous):

what I would do is flip the rows that Hero wrote because then you would already have a 1 in the upper left hand corner which is what we need.

OpenStudy (anonymous):

flip them in which way

OpenStudy (anonymous):

1 -1 | 3.5 2 5 | 14

hero (hero):

[1 -1 3.5] [2 5 14]

OpenStudy (anonymous):

1 -1 | 3.5 2 5 | 14 add (-2 times row 1) to row 2 I did this because we want a 0 in the bottom left hand corner. A couple matrix rules you should know. 1) you can add and subtract rows 2) you can rearrange rows 3) you can multiply a row by a scaler 4) you CANNOT multiply rows

OpenStudy (anonymous):

1 -1 | 3.5 2 5 | 14 add (-2 times row 1) to row 2 New Matrix (Row 1 stays the same, we assume that after every row operation we return that row to the original) 1 -1 | 3.5 0 7 | 7

OpenStudy (anonymous):

why add -2 in particular?

OpenStudy (anonymous):

because 1*-2 = -2 and when we add that to row 2 we get a zero in the bottom right 2+-2=0

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

Now we want the bottom right hand corner to be a 1. We can just divide this row by 7. Remember I said that we are allowed to multply a row by a scaler, in this case 1/7. 1 -1 | 3.5 0 7 | 7 divide by 7 New Matrix 1 -1 | 3.5 0 1 | 1

OpenStudy (anonymous):

Now we want the upper right hand corner to be a 0. so we can just add row 2 to row 1 1 -1 | 3.5 add row 2 to row 1 0 1 | 1 1 0 | 4.5 0 1 | 1 We have 1's in our diagonal. Everything else a 0. We are done. What does this mean?

OpenStudy (anonymous):

1 0 | 4.5 0 1 | 1 The first column if you remember is our x's and the second is our y's. Row 1 is an equation as well as row 2. (1)x + (0)y = 4.5 (0)x + (1)y = 1 x=4.5 y=1 done

OpenStudy (anonymous):

any questions?

OpenStudy (anonymous):

how do you know when to divide, subtract or add? or is there a particular rule in which order it goes?

OpenStudy (anonymous):

can you use any method as long as you get the result wanted?

hero (hero):

It an intuitive thing

OpenStudy (anonymous):

You want to work down the rows one at a time getting the necessary numbers. If I had to put it in steps I'd say this: 1) Get the top left corner a 1 first by dividing this row by whatever is necessary 2) Get the remaining numbers in the first column all zero by multiplying row 1 by the necessary scaler and adding that to the desired row 3) Move to the second row and get the 2 number to be a 1 by multiplying it by a scaler 4) Get the remaining numbers beneath that a zer by multiplying row 2 by the necessary number and adding that to the desired row 5) repeat these steps til you have gotten to the last 1 in the bottom right. 6) Now you need to work back up. making the numbers in the row above zeros. It is the same process but in reverse

OpenStudy (anonymous):

Hope that helps. I'm getting off. Good luck. It takes practice to do them quickly

OpenStudy (anonymous):

Thank you sooo much!!!